OpenStudy (anonymous):

Let f(x,y)=(2x−y)^7. Then ∂^2f/∂x∂y = ∂^3f/∂x∂y∂x= ∂^3f/∂x^2∂y =

OpenStudy (psymon):

Im not positive of the notation, new stuff to me, but im guessing its saying partial with respect to x first then to y for the first one. Just to make sure Im getting what its asking.

OpenStudy (anonymous):

\[ \frac{ ∂^{2}f }{ ∂x∂y} = ? \] \[ \frac{ ∂^{3}f }{ ∂x∂y∂x} = ? \] \[ \frac{ ∂^{3}f }{ ∂x^{2}∂y} = ? \]

OpenStudy (anonymous):

@Psymon I think its asking the 2nd partial for the first one/

OpenStudy (psymon):

Yeah, 2nd parital, but with the first one being with respect to x and then the second being with respect to y. Of course I could be wrong, Either way, partial with respect to x first so we can get the idea. So, if we differentiate with respect to x, we treat y as a constant. In fact, if it makes it less confusing to do, you could replace y with a constant. If I differentiated the function and pretended y was the number 1, I could do something liek this: \[(2x-1)^{7} = 7(2x-1)^{6}(2) = 14(2x-y)^{6}\] Not that making it a 1 is necessary, but maybe its a step that makes it less confusing. Now if I do the derivative of this result with respect to y, I pretend x is a constant. SO in the same manner, possibly to avoid confusion, I can replace 2x with 1 to prevnt it from confusing me \[14(1-y)^{6} = 84(1-y)^{5}(-1) = -84(2x-y)^{5}\]

OpenStudy (anonymous):

so for the next one we take −84(2x−y)^5 and do another partial in respect to x?

OpenStudy (psymon):

Right

OpenStudy (anonymous):

so -840 (y-2 x)^4. No what happens on the 3rd one?

OpenStudy (anonymous):

now*

OpenStudy (anonymous):

or is that the same question?

OpenStudy (psymon):

So the 3rd one we have to backtrack and instead of doing partial to y for the 2nd derivative we do it to x a second time. So the first partial was \[14(2x-y)^{6}\]Now we just do partial again with respect to x \[84(2x-y)^{5}(2) = 168(2x-y)^{5}\]Now finally we do the 3rd derivative but with respect to y this time\[840(2x-y)^{4}(-1) = -840(2x-y)^{4}\]

OpenStudy (anonymous):

came to same answer haha. Thanks!

OpenStudy (psymon):

Yep, np : )